OpenStudy (anonymous):

A triangle drawn on a map has sides of lengths 8.0 cm, 10 cm, and 15 cm. The shortest of the corresponding real-life distances is 110 km. Find the longest of the real-life distances. Round your answer to the nearest whole number.

OpenStudy (agent47):

8/110 = 15/x x=110*15/8 = 206.25 and you need to brush up on similar triangles.

OpenStudy (anonymous):

thanks